OpenStudy (anonymous):

Find tan θ if sec θ = (sqrt26)/5 and sin θ < 0.

OpenStudy (anonymous):

\[\tan\theta = \frac{\sin\theta}{\cos\theta}\] \[\sec\theta = \frac{1}{\cos\theta}\]

OpenStudy (anonymous):

I got cosθ= (5sqrt26)/26

OpenStudy (anonymous):

also you can form a triangle from the secant ratio since: \[\sec\theta = \frac{hyp}{adj}\] 

OpenStudy (anonymous):

since the tan requires the opposite side: \[\tan\theta = \frac{opp}{adj}\] you should solve missing side with pythagorean

OpenStudy (anonymous):

So tanθ= 1/5 ?

OpenStudy (anonymous):

right.

OpenStudy (anonymous):

do the other questions your working on ask you to find an angle? if so i think that's why they give you: sin θ < 0

OpenStudy (anonymous):

but in this context it appears to be extra information.

OpenStudy (anonymous):

Well 1/5 isn't an option as an answer, but - 1/5 is...

OpenStudy (anonymous):

@Mathmuse

jimthompson5910 (jim_thompson5910):

since sin(theta) < 0, cos(theta) > 0 and tan = sin/cos, this means that tan(theta) is negative

OpenStudy (anonymous):

Thank you @jim_thompson5910 !

jimthompson5910 (jim_thompson5910):

np